OpenStudy (anonymous):

Medals given! A system of equations is shown below: 6x = -7y + 11 5x + 2y = 13 What is the solution to this system of equations?

OpenStudy (***[isuru]***):

6x = -7y + 11 6x + 7y = 11 ---(1) (1) x 2 12 x + 14 y = 22---(2) 5x + 2y = 13 ---(3) (3) x 7 35x + 14y = 91 ---(4) now.. subtract (2) from (4) (4) - (2) = ? can u continue it now ?

OpenStudy (anonymous):

I think so. Is it 3, -1?